OpenStudy (anonymous):

Find the slope of the line containing the points (1, -4) and (-1, 3) , -5/4??????

OpenStudy (anonymous):

[3-(-4)]/(-1-1) = -7/2

OpenStudy (anonymous):

i got -2/7 ?

OpenStudy (anonymous):

yep the equation is m= slope m=(y_2_-Y_1_/x_2_-X_1_)

OpenStudy (anonymous):

so y2=3 y1=-4 x2 =-1 and x1=1

OpenStudy (anonymous):

so, ((3-(-4))/(-1-1)))

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

so 7/-2 ?

OpenStudy (anonymous):

ya and that negative will make the whole term negative leaving you with -7/2

OpenStudy (anonymous):

its really simple if you can remember the equation i gave you at first m=(y2-y1)/(x2-x1)